A square field had 5 meters added to its length and 2meters added to its width.The field had an area of 130 cm .Find the length of the original field.
Found 2 solutions by ankor@dixie-net.com, ikleyn:
Answer by ankor@dixie-net.com(22740)

A square field had 5 meters added to its length and 2meters added to its width.
The field had an area of 130 sq/m .
Find the length of the original field.
:
s = side of the original square
(s + 5) = the length of the new field
(s + 2) = the width
:
(s+5)*(s+2) = 130
FOIL
s^2 + 2s + 5s + 10 = 130
s^2 + 7s + 10 - 130 = 0
s^2 + 7s - 120 = 0
Factors to
(s+15)(s-8) = 0
A positive solution
s = 8 meters is the side of the original square
;
:
Check: (8+5)*(8+2) = 130

Answer by ikleyn(52878)